OpenStudy (anonymous):

Does the data in the table represent a direct or inverse variation? Write an equation to model the data in the table. A) direct variation; y = 12/x B) direct variation; y = 12x C) inverse variation; xy = 12 D) inverse variation; y/x = 12 ______________________ X | 2 | 4 | 8 | 12 Y | 6 | 3 | 3/2 | 1

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

what do you get when you multiply each pair?

OpenStudy (anonymous):

\[2\times 6\\ 4\times 3\\ 8\times \frac{3}{2}\\ 12\times 1\] they are all the same